OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU METAL AND FAN please help!!! Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (campbell_st):

well its really saying z + 0.3 = -2.4 so to find z subtract 0.3 from both sides of the equation z + 0.3 = -2.4 - 0.3 -0.3 -------------- z = ? what is the value of ?

OpenStudy (winsteria):

-2.1??? @campbell_st

OpenStudy (campbell_st):

no... you need to subtract... to get -2.1 you would have added 0.3 to -2.4

OpenStudy (winsteria):

Uh idk what it is ;-;

OpenStudy (campbell_st):

well if you were on the numberline at -2.4 subtracting 0.3 means going left... here an online calculator to use http://web2.0calc.com/

OpenStudy (winsteria):

Is it -2.7 or 2.1 or...

OpenStudy (campbell_st):

it's the 1st one

OpenStudy (winsteria):

Okay so -2.7 thank you so much! :)